OpenStudy (anonymous):

For each pair of functions, find f(g(x)) and g(f(x)). 1. f(x) = x-3/2, g(x) = 2x-3 2. f(x) = x+5/2, g(x) = x^2

OpenStudy (anonymous):

1. f(x) = x-3/2, g(x) = 2x-3 f(g(x)) = g(x)-3/2 f(g(x)) = 2x-3-3/2 does that help?

OpenStudy (anonymous):

Yes and this is what I did for the first problem: = f(2x-3) = (2x-3) -3 = -6x+9/2 Can I leave the answer to be -6x+9/2?

OpenStudy (anonymous):

I'm stuck on the first problem with = g(x-3/2) = 2(x-3/2) -3

OpenStudy (anonymous):

how did you get -6x?

OpenStudy (anonymous):

Oh wait I distributed -3 and I see that you can only subtract it... My bad. So would it be like this then: = f(2x-3) = (2x-3) - 3/2 = 2x-6/2 = 2x - 3 ?

OpenStudy (anonymous):

= f(2x-3) = (2x-3) - 3/2 = 2x-3 - 3/2 .................. (3=3/1=6/2) = 2x-6/2 - 3/2 = 2x - 9/2 now try g(f(x))

OpenStudy (anonymous):

= g(x-3/2) = 2(x-3/2)-3 = (x-3)-3 = x-6

OpenStudy (anonymous):

= g(x-3/2) = 2(x-3/2)-3 = (x-3)-3 ← why didn't you distribute the 2 to the x? = x-6

OpenStudy (anonymous):

I missed that one -__- . Gah I'm so bad at math. So: = g(x-3/2) = 2(x-3/2)-3 = (2x-6) - 3 = 2x-9

OpenStudy (anonymous):